use double angle formulas to verify the identity algebraically.
6sinxcosx = 3sin2x

Answer by fractalier(6550)

Okay, sin 2x = sin(x + x) = sin x cos x + sin x cos x = 2 sin x cos x
3 sin 2x = 3(2 sin x cos x) = 6 sin x cos x